Alexis draws 300 circles with diameter 1, Brianna draws 400 squares with side length 1, and Cyril draws 500 equilateral triangles with side length 1. If they all draw with the same type of pen, who uses the most ink?
Answer by Edwin McCravy(8909)


Alexis draws 300 circles with diameter 1, 
Brianna draws 400 squares with side length 1, 
and Cyril draws 500 equilateral triangles with 
side length 1. If they all draw with the same 
type of pen, who uses the most ink?

When you draw a circle you only draw the 
circumference, unless you shade it in, and I
assume they didn't.  When you draw a square or a 
triangle, and don't shade it in, you only draw 
the perimeter. (If they shaded their figures
we'd have to use the areas instead).

A circle of diameter d=1 has a circumference
C = pd = p(1) = p.  

So the sum of all the circumferences of Alexis' 
300 circles is 300p.  That's about 942.4777961.


A square with side s=1 has a perimeter 
P = 4s = 4(1) = 4

So the sum of all the perimeters of Briannas' 400 
squares is 400(4) = 1600.


An equilateral triangle with side s=1 has a 
perimeter P = 3s = 3(1) = 3

So the sum of all the perimeters of Cyril's 500 
equilateral triangles is 500(3) = 1500.

So Brianna uses the most ink with her squares.


[If they had shaded them in Cyril would have 
used the most ink).

Edwin